# LOCUS – LOcalization with Channel Uncertainty and Sporadic Energy

### v2

@@ 6. Results
LOCUS with
SALSA
20,560M 178B 1.19s 132.76J
6.1. Evaluation on Different Sound Source Localization Algorithms
Table 11 provides a complexity and overhead analysis for SALSA and SELDNet, both with and without LOCUS. It highlights that SALSA has 34,093.87×34,093.87$\times$34 , 093.87 × more parameters and requires 8.45×8.45$\times$8.45 × more floating-point operations per second (FLOPS) compared to SELDNet. Figure 15 shows that, when combined with LOCUS, the simpler SELDNet (with only $595K$ parameters) achieves an $E_{DoA}$ that is $14.01$ degrees lower than SALSA. Moreover, LOCUS introduces only a minimal runtime overhead of $3.45-5.19\%$ and a slight increase in power consumption of $7.12-7.33\%$, which remains lower than SALSA alone.
@@ 6.2. Comparison with Baseline Algorithms
On the contrary, LOCUS uses the information estimation, recovers the missing features, and reduces $E_{DoA}$ to $36.90$ and $28.62$ degree. Which is a $13.95-76.72\%$ improvement over all the baselines on DCASE. LOCUS also outperforms all the compared baselines $32.55-67.04\%$ on LargeSet.
@@ 6.3. Effect of Missing Data Percentage
Figure 12. Performance of SELDNet, $RTrain_{p}$ and LOCUS trained with $p\%$ MDP.
@@ 6.4. Ablation Study
This section evaluates the contribution of each component of LOCUS using the DCASE dataset and SELDNet SSL network.
@@
6.5. Beyond Audio
To validate LOCUS across domains, we test it on the WiFi-based user localization algorithm DLoc (Ayyalasomayajula et al., 2020) using a large-scale WiFi CSI dataset with 16 channels (4 access points, 4 antennas each) and 75% missing data (MDP=75%). The dataset covers two indoor environments spanning 2000 sq. ft. across 8 scenarios.
Localization Network. We follow the two-step image-based feature extraction process proposed by the authors. The model architecture includes a single encoder and two decoders (consistency and location). The encoder processes input heatmaps from all access points to generate a concise representation, which is then fed into the decoders.
Result. Figure 15 shows the baseline DLoc with a 90th percentile error of 2.98 meters, while retraining ($DLoc_{retrain}$) increases the error to 3.29 meters due to varied missing data scenarios. In contrast, LOCUS reduces the error to 2.13 meters, achieving a 35.25% gain over $DLoc_{retrain}$ and 25.52% over the baseline.
@@ 7. Real World Evaluation
Figure 16. Data corruption example with energy harvester.
@@
Figure 17. Comparison of LOCUS with SELDNet, and $RTrain_{75}$ with solar and radio frequency (RF) harvester.
This section presents our real-world evaluation of LOCUS across three environments under two energy scenarios, where the microphone array is powered by ambient energy.
Table 2. Environment Description.
{tblr}
width = colspec = Q[38]Q[350]Q[835], rowodd = Mercury, No & Dimension ($ft$) Description
#1 $23.5\times 46\times 10$ A large room with no window and lots of furniture and electrical equipment.
Energy Harvesting Setup. We employ an Ethylene Tetrafluoroethylene (ETFE) solar panel (sol, 2020) with a step-up regulator (ltc, 2015) and a Powercast harvester-transmitter system (pow, 2016a, b) to harvest radio frequency (RF) and solar energy across three environments. The solar panel is placed near a window, with intermittence caused by clouds and human shadows, while the RF transmitter and receiver are positioned on opposite walls of a corridor, experiencing interruptions from uncontrolled human traffic.
Acoustic Data Collection Setup. We utilize two MAX78000 microcontrollers to capture data from a four-channel microphone array (202, 2023), each powered by independent energy harvesters, contributing to intermittent missing channels. Since each microcontroller has only one I2S interface, two are needed to handle the array’s four-channel data. A speaker is positioned at 18 locations in 10-degree increments, with varying distances between the speaker and the array across three different environments, as detailed in Table 2. Precise measurements of the speaker’s position relative to the microphones serve as ground truth.
We collect a polyphonic dataset featuring 1-3 concurrent audio classes, using 96 one-minute segments from seven audio classes in the NIGENS database (Trowitzsch et al., 2019). The environments introduce uncontrolled noises such as HVAC sounds, conversations, and other ambient background noise. A total of 96 minutes of audio is gathered, with 72 minutes used for training and 24 minutes for evaluation. Figure 17 shows example data with a solar panel as the energy harvester.
@@
Results. Figure 17 demonstrates that LOCUS significantly outperforms both pre-trained and retrained SELDNet, reducing $E_{DoA}$ by $25.87-60.55\%$ and $29.65-59.46\%$, respectively, in solar harvester scenarios. Interestingly, retrained SELDNet shows higher $E_{DoA}$ than the pre-trained model in environments 1 and 3. This occurs because retraining is done with $MDP=75$%, whereas the real-world introduces stochastic variations in missing information. Despite being trained on the same $MDP=75$% data, LOCUS effectively recovers stochastically missing information, highlighting the robustness of the algorithm.
@@
In the RF harvester scenario, where sharper and more frequent energy variations occur, the challenge is more dynamic compared to solar. Figure 17b shows that RF harvesting results in more missing channels, leading to higher errors for the pre-trained SELDNet across all environments. Pre-trained and retrained SELDNet experience $118-157\%$ and $33-160\%$ higher $E_{DoA}$ than LOCUS. In contrast, LOCUS shows only $3.84-22.50\%$ more error compared to the scenario with no missing channels, further demonstrating its generalizability and robustness. Notably, in contrast to the solar harvester, retrained SELDNet has lower error than the pre-trained model in environment 1, suggesting that retraining on a fixed MDP does not adapt well to lower MDP conditions, making it less suitable for real-world applications.
@@ 8. Related Work
Intermittent System. Initial research in intermittent computing emphasized maintaining data consistency and integrity under unreliable power conditions. Systems such as Mementos (Ransford et al., 2011) introduced checkpointing strategies to periodically store system states, thus mitigating data and computation loss due to power interruptions. Although checkpointing reduces data loss, it does not fully address the issue of missing data caused by power interruptions between checkpoint intervals. Recent systems, such as Greentooth (Babatunde et al., 2024), propose resilient architectures for energy-harvesting devices to better withstand power outages. Other recent works, including FreeML (Farina et al., 2024) and SONIC (Gobieski et al., 2018), explore efficient inference of deep neural network (DNN) models on intermittent systems. In addition, (Majid et al., 2020b) demonstrate audio sensing capabilities on intermittently-powered devices; however, their system only captures audio samples of short events ( 283 ms), which remains insufficient for many real-world applications requiring extended and continuous data collection.
Missing Information Retrieval. To recover missing information from sensors, statistical methods (Jerez et al., 2010) like mean imputation (Allison, 2001), hot deck (Little and Rubin, 2019), and multiple imputations (Rubin, 1996) are often used. However, these methods focus on single-channel data and fail for long missing data sequences (Hasan et al., 2021).
In the deep learning world, missing information retrieval from a single sensor has been studied by the Natural Language Processing (NLP) and Computer Vision (CV) Community. While works in the NLP domain focus on predicting missing words and syllables in texts (Katwe et al., 2020), while CV works focus on restoring old images(Zhang et al., 2017), removing obstacles(Christiansen et al., 2016), or eliminating raindrops (Peng et al., 2020). Some other works focus on learning the input distribution (Alemi et al., 2016) to denoise the noisy inputs but can not handle the missing information. The most related work to acoustic missing information recovery is SoundSieve (Monjur et al., 2023), which recovers missing segments from single-channel audio that occurred due to energy intermittence for a sound event classification task. However, this method does not consider multiple channels and takes phase information into account for imputation. Few works on speech enhancements have looked into recovering corrupted multi-channel data (Taherian et al., 2022). However, unlike source localization, speech enhancement only partially depends on hard-to-recover spatial information and thus can not be directly used for localization tasks. Though some recent works focus on recovering missing modalities in multimodal sensing (Ma et al., 2022, 2021), no work focuses on interrelated sensors of a single modality.
@@
Sound Source Localization. Classic sound source localization (SSL) methods, e.g., MUSIC (Gupta and Kar, 2015), independent component analysis (Noohi et al., 2013), and sparse models (Yang et al., 2018), perform poorly for under-determined scenarios. Though DNN has shown promising results for SSL (Ferguson et al., 2018; Yiwere and Rhee, 2017; Hirvonen, 2015), none of the existing works focus on missing or corrupted information.
@@
Attention Mechanism. Attention learns to focus on essential information and ignore the rest selectively (Colombini et al., 2014). For multi-channel input, attention explores the channel characteristics and estimates the channel state information (Gao et al., 2021). Though commonly used later in a network, early attention improves performance with substantial margins (Hajavi and Etemad, 2020) motivating our InFo component.
@@ 9. Discussions and Limitations
This section describes the design choices and limitations and future directions of LOCUS.
Figure 18. Effect of retraining the downstream task.
Impact of Joint Training. LOCUS requires joint training of InFo, LaFS, and the sound source localization network to effectively recover missing information. To assess the need for retraining, we tested a setup where InFo and LaFS were added to a pre-trained, frozen SELDNet and compared it to LOCUS with a jointly trained localization network. Figure 18 shows that the pre-trained network increases $E_{DoA}$ by $14.3-103.08\%$ compared to no missing channels, while LOCUS reduces $E_{DoA}$ by $1.9-16.62\%$. This is because InFo alters the feature distribution, requiring the localization network to adapt during training, making retraining essential. Future work will explore domain generalization to potentially eliminate the need for retraining and make LOCUS a plug-and-play solution for missing information recovery.
@@
Generative Models for Information Retrieval.. Generative models like VAEs and GANs have shown promise in recovering information, but VAEs often struggle with reconstruction accuracy and diversity due to their simplified assumptions, making them inadequate for complex, discrete data. GANs, while powerful, require large datasets and are sensitive to hyper-parameters, making them difficult to train in dynamic environments. While these models could replace the auto-encoder in LaFS, we find that a simpler, easier-to-train auto-encoder performs comparably with less overhead, making it the better choice for multi-channel information retrieval in LOCUS.
LOCUS for Multi-Sensor Systems. Multi-modal systems involve data from different sources, such as video, audio, and text, which may or may not be interrelated. In contrast, multi-sensor data, like a 12-lead ECG, collects interrelated data from multiple channels. This paper demonstrates LOCUS’s effectiveness in a multi-sensor setup and suggests its potential expansion into multi-modal domains. For example, in video conferencing, LOCUS could retrieve corrupted audio data from video and text to enhance overall quality.
Microphone Geometry Dependency. LOCUS operates on top of existing sound source localization methods and, like them, relies on the microphone-array geometry through time difference of arrival (TDoA) and triangulation. As a result, it performs poorly when used with a different geometry, requiring retraining each time the geometry changes. For example, the network trained in section 6.2 cannot be applied directly in section 7. Our future work plans to explore domain generalization techniques to address this dependency.
@@
Large Real-World Localization Dataset. In the absence of large public datasets for sound source localization, we created and released a 50-hour simulated dataset. While useful for training DNNs, simulated data may lack the diversity and complexity of real-world conditions due to inherent assumptions. Future work will focus on developing a comprehensive real-world sound source localization dataset.
Image Recovery Limitations for Acoustic Features. While deep learning methods for image pixel recovery are effective (Zhang et al., 2017; Chen et al., 2022), they may not work for frequency domain features like GCC and MFCC. Unlike images, where neighboring pixels often share similar values, these acoustic features lack such similarity, making image recovery methods unsuitable for restoring missing acoustic channel information.
10. Conclusion
In this paper, we develop an information retrieval-based sound source localization system that can hold its performance in the presence of missing or corrupted data. We leverage the interdependence among different random variables among the channels to recover missing information. Meanwhile, classical algorithms to handle missing data points are insufficient due to their limitations, e.g., the requirement of prior knowledge and performance degradation on high and complex data distributions. LOCUS overcomes these challenges by proposing InFo, responsible to quantify how much and where the corruptions are and LaFS to interpolate them. Our evaluation shows that our proposed LOCUS can achieve up to $36.91\%$ less degree of arrival error than just retraining the localization model using data with missing channels on DCASE dataset. Our real-world evaluation shows $25.87-59.46\%$ performance gain in stochastically missing information.

### v3

@@ 6.1. Evaluation on Different Sound Source Localization Algorithms
LOCUS- SALSA 20,560M 178B 1.19s 132.76J
Table 2 provides a complexity and overhead analysis for SALSA and SELDNet, both with and without LOCUS. It highlights that SALSA has 34093.87×34093.87$\times$34093.87 × more parameters and requires 8.45×8.45$\times$8.45 × more floating-point operations per second (FLOPS) compared to SELDNet. Figure 13 shows that, when combined with LOCUS, the simpler SELDNet (with only $595K$ parameters) achieves an $E_{DoA}$ that is $14.01$ degrees lower than SALSA. Moreover, LOCUS introduces only a minimal runtime overhead of $3.45-5.19\%$ and a slight increase in power consumption of $7.12-7.33\%$, which remains lower than SALSA alone.
@@ 6.2. Comparison with Baseline Algorithms
On the contrary, LOCUS, recovers the missing features, and reduces $E_{DoA}$ to $36.90$ and $28.62$ degree. Which is a $13.95-76.72\%$ improvement over all the baselines on DCASE. LOCUS also outperforms all the compared baselines $32.55-67.04\%$ on LargeSet.
@@ 6.4. Ablation Study
This section evaluates the contribution of each component of LOCUS using the DCASE dataset and SELDNet SSL network. Our ablation shows that this soft fusion using $I(\tilde{F})$ outperforms alternatives such as direct interpolation or naïve masking.
@@
Table 3. Environment Description for Real World Evaluation.
{tblr}
width = colspec = Q[38]—Q[350]—Q[835], rowodd = Mercury, [1.5pt] No Dimension ($ft$) Description
[1pt] #1 $23.5\times 46\times 10$ A large room with no window and lots of furniture and electrical equipment.
[1.5pt]
@@ 7. Real World Evaluation
This section evaluates LOCUS in three real-world environments with two energy scenarios, using ambient energy to power the microphone array.
Energy Harvesting Setup. We use an ETFE solar panel with a step-up regulator and Powercast harvester-transmitter system to harvest RF and solar energy across three environments. The solar panel, placed near a window, experiences intermittence from clouds and shadows, while the RF system faces disruptions from human traffic. Energy traces are collected from multiple MCU locations, each powered by different harvesters.
@@
Figure 15. LOCUS with SELDNet recovers missing features more effectively and achieves lower $E_{DoA}$ than $RTrain_{75}$ SELDNet for solar and RF harvested devices.
Acoustic Data Collection Setup. We use two MAX78000 microcontrollers to record from a four-channel microphone array, with each MCU connected to two microphones and powered by independent RF or solar harvesters. This distributed setup simulates real-world conditions where environmental factors, such as RF shadowing, multipath, or occlusion, lead to asynchronous power availability across nodes, resulting in partial channel dropouts. Each MCU logs data locally, and synchronization is achieved using periodic sync tones emitted by the sound source and captured by all microphones. This approach allows for offline alignment of data without the need for real-time synchronization. Due to I2S constraints, we employ two MCUs to manage the microphone array, ensuring independent and efficient data collection across the system.
We collect a polyphonic dataset with 1–3 concurrent audio classes, using 96 one-minute segments from seven classes in the NIGENS database (Trowitzsch et al., 2019). The speaker is placed at 18 azimuthal angles (10° increments) across three environments (Table 6.4). Precise speaker locations are measured as ground truth, and background noise introduces natural variability. A total of 96 minutes of data are collected, with 72 for training and 24 for evaluation.
@@
Results. Figure 15 shows that LOCUS outperforms both pre-trained and retrained SELDNet, reducing $E_{DoA}$ by $25.87-60.55\%$ and $29.65-59.46\%$ in solar scenarios. Retrained SELDNet has higher $E_{DoA}$ than the pre-trained model in environments 1 and 3 due to retraining on $MDP=75$% data, which does not account for stochastic variations. In contrast, LOCUS effectively recovers missing information.
@@
In the RF harvester scenario, where energy fluctuations are more frequent, Figure 14b shows that RF harvesting leads to more missing channels, causing higher errors for pre-trained SELDNet. Pre-trained and retrained SELDNet show $118-157\%$ and $33-160\%$ higher $E_{DoA}$ than LOCUS. In contrast, LOCUS only shows $3.84-22.50\%$ more error than the no-missing-channels scenario, demonstrating its robustness. In contrast to solar harvesting, retrained SELDNet performs better than the pre-trained model in environment 1, suggesting that retraining on a fixed MDP does not adapt well to lower MDP conditions.
@@ 8. Related Work
Intermittent System. Early research in intermittent computing focused on data consistency under unreliable power, using checkpointing strategies to mitigate data loss during power interruptions. Systems like Mementos (Ransford et al., 2011) implemented checkpointing but failed to address gaps in data during periods between checkpoints. Recent systems such as Greentooth (Babatunde et al., 2024) and FreeML (Farina et al., 2024) explore resilient architectures for energy-harvesting devices, while SONIC (Gobieski et al., 2018) improves DNN inference under intermittent power. Additionally, (Majid et al., 2020b) shows audio sensing on intermittently-powered devices but is limited to short events ( 283 ms), insufficient for continuous data collected in real-world applications.
Missing Information Retrieval. Statistical methods such as mean imputation (Allison, 2001), hot deck (Little and Rubin, 2019), and multiple imputations (Rubin, 1996) are used to handle missing data, but these methods are mainly designed for single-channel data and struggle with long sequences of missing data (Hasan et al., 2021). In the deep learning domain, NLP (Katwe et al., 2020) and CV (Zhang et al., 2017) have explored missing data recovery, but these approaches mainly focus on text or image restoration. The closest work to our approach is SoundSieve (Monjur et al., 2023), which addresses missing segments in single-channel audio due to energy intermittence, but it does not consider multi-channel systems. Other speech enhancement works, such as (Taherian et al., 2022; Mohapatra et al., 2024), focus on recovering corrupted multi-channel data, but these are not directly applicable to sound source localization, which requires spatial data recovery. Recent multimodal work (Ma et al., 2022, 2021) recover missing modalities, but ignores interrelated sensors in a single modality.
@@
Sound Source Localization. Classic SSL methods, like MUSIC (Gupta and Kar, 2015), independent component analysis (Noohi et al., 2013), and sparse models (Yang et al., 2018), struggle with under-determined scenarios. Deep learning-based SSL has shown promise (Ferguson et al., 2018; Yiwere and Rhee, 2017), but fails in presence of missing or corrupted data.
@@
Attention Mechanism. Attention mechanisms focus on critical information while ignoring irrelevant data (Colombini et al., 2014). In multi-channel input, attention is used to explore channel characteristics and estimate channel state information (Gao et al., 2021). Although typically applied later in a network, early attention has shown substantial performance improvements (Hajavi and Etemad, 2020), motivating the use of InFo in our approach.
@@ 9. Discussions and Limitations
Figure 16. Result of WiFi localization at $MDP=75\%$.
Beyond Audio and DoA. While LOCUS is designed for sound source localization, its feature recovery formulation is modality-agnostic. To demonstrate generalizability, we apply it to Wi-Fi CSI-based localization (distance estimation) using amplitude maps analogous to MFCCs. We evaluate LOCUS on the DLoc model (Ayyalasomayajula et al., 2020) with a large-scale dataset covering two indoor environments and 16 CSI channels (4 access points × 4 antennas) under 75% missing data (MDP = 75%). We adopt DLoc’s two-step architecture: an encoder processes CSI heatmaps, and two decoders predict spatial consistency and location. As shown in Figure 16, LOCUS reduces the 90th percentile localization error to 2.13 m 35.25% gain over retrained DLoc (3.29 m) and 25.52% over the original (2.98 m). This demonstrates LOCUS ’s applicability beyond audio, including in RF sensing with incomplete data.
Offline Processing. LOCUS is designed for post-hoc analysis, where data is logged during deployment and processed offline, making it suitable for batteryless or energy-constrained systems where real-time inference is impractical. This allows us to employ complex recovery mechanisms without on-device computation costs, though it limits its use in ultra-low-latency applications. Future work could explore lightweight, on-device approximations and integrate recovery-aware sensing strategies. Additionally, while we use sync tones for alignment, drift-aware synchronization could be explored for long-term distributed deployments.
Impact of Joint Training. LOCUS is jointly trained with the localization network to address feature distribution shifts caused by InFo and LaFS. As shown in Figure 17, joint training improves performance significantly. The recovery blocks are modular and can be trained separately for integration into existing systems, with future work exploring generalization strategies to reduce retraining requirements.
Figure 17. Impact of joint training on $E_{DoA}$.
@@
Generative Models for Information Retrieval. Generative models like VAEs and GANs are powerful but may struggle with reconstruction accuracy, particularly for discrete data. While they could replace the auto-encoder in LaFS, we find that a simpler auto-encoder performs comparably with less training overhead, making it a more efficient choice for multi-channel information retrieval in LOCUS.
LOCUS for Multi-Sensor Systems. While this paper focuses on multi-sensor systems like microphone arrays, LOCUS could be extended to multimodal systems (e.g., video, audio, text). For instance, LOCUS could recover missing audio data from video or text sources in a video conferencing setup, improving the overall quality.
Microphone Geometry Dependency. LOCUS relies on microphone-array geometry for localization and functions as a preprocessing layer that reconstructs reliable features before they are passed to downstream SSL architectures. This design allows us to support existing geometry-dependent methods without requiring modifications to their core structure. Future work will explore domain generalization techniques to address this limitation.
@@
Large Real-World Localization Dataset. We created and released a 50-hour simulated dataset to train DNNs for sound source localization. However, simulated data may lack the complexity of real-world environments, and future work will focus on developing a comprehensive real-world dataset for sound source localization.
MDP incorporating number of missing channels. While incorporating the number of missing channels could add granularity, we excluded it since MDP is solely an evaluation metric, not used during training. Using a consistent MDP across varying missing-channel scenarios introduces a more challenging evaluation, better showcasing our method’s robustness; jointly modeling temporal extent and channel count remains a promising direction for future work.
Image Recovery Limitations for Acoustic Features. While image pixel recovery methods are effective for images, they are unsuitable for frequency-domain features like GCC and MFCC, which do not exhibit the same neighbor similarity. Thus, we cannot directly apply image recovery methods to restore missing acoustic channel data.
Handling Sparse Inputs. While LOCUS is not designed to operate on sparse inputs, it reconstructs missing features to restore dense representations before localization. Future work could explore architectures operating on sparse or partially observed features to reduce reconstruction overhead.
System-Level Exploration. A full exploration of communication overhead, runtime efficiency, and hardware acceleration is beyond the scope of this work. However, future efforts could focus on lightweight on-device approximations and energy-aware communication strategies for embedding LOCUS in fully deployed systems. To support on-device adaptation, we plan to explore a more comprehensive set of baselines and adopt efficient model architectures tailored for resource-constrained environments. Additionally, we plan to reduce feature dimensionality and apply model compression techniques such as pruning and quantization (Biswas et al., 2025) to minimize memory and computational demands.
In this paper, we present LOCUS, an information retrieval-based sound source localization system designed to perform effectively despite missing or corrupted data. By leveraging the interdependence of data across channels, LOCUS recovers missing information that traditional algorithms, which require prior knowledge and struggle with high-complexity data, cannot handle. The InFo component quantifies both how much and where the data is corrupted, while LaFS interpolates the missing features. Our experiments demonstrate that LOCUS achieves up to 36.91% lower degree of arrival (DoA) error compared to retrained models using missing-channel data in the DCASE dataset. Additionally, in real-world scenarios with stochastically missing information, LOCUS shows performance gains of 25.87-59.46%, highlighting its robustness and effectiveness in real-world conditions.
Acknowledgment
This research was supported by funding from the NSF CNS-2347692. We gratefully acknowledge their support in enabling this work.